Which of the following languages are regular. Prove (by providing a regular expression) or disprove.

a. L1 = {ai bj ck dl | (i + j)mod 2 = (k + l)mod 2 , i, j, k, l ≥ 0}

b. L2 = {ai bj ck dl | (i + j) = (k + l), i, j, k, l ≥ 0}
Answer #1

1. a, b, c, d * prove above language a) La & a bicka (Citj) mod2 = (KHL) mod2, i, j, k, 120} are infinitely produced strings.Care & Check xyk z el for different values of K. Kal =) ) a call bred a aa bccd El a can bred K=2 =) aaabeed &L Reason - i=3,in 2, j = 1, K22 dal a str=a²b²c²d zaabed where strela Stelp!- Nowo split the string into 3 parts. a albccd al y z da.y=0,2 =

itj= 3+1 = 4 KAL - 2+1=3 (i+) + (K+1) Caselii) fail. conclusion - since Caselii) fail our assumption that language L2 as regu